# Adaptive behavior with stable synapses

### v1

@@ Results
Figure 2: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of 10 units (described by the vector $\bm{z}(t)$, depicted in pink) is utilized for learning a periodic trajectory, following the prescription of reservoir computing,. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios.
@@
We aim to demonstrate that such an approach can be generalized to non-stationary scenarios.
We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm, we employ an RNN (Fig. 2A, pink), whose dynamics follows Equation (7), to extract temporal features $\bm{z}(t)$. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 2B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN. The readout virtual weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
@@
Here, the network $W_{\Theta^{w}}$$({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)=\textrm{GMNN}_{\Theta^{w}}($$\bm{z}$$,e|\gamma_{e}e)$ is dedicated to estimating the gradient of virtual weights, while $Y_{\Theta^{y}}$(((($\bm{z}$$,w)=\textrm{GMNN}_{\Theta^{y}}($$\bm{z}$$,w|\gamma w)$ is tasked with estimating the predicted $y(t)$ as a function of the new virtual weights (Fig.2A green and orange respectively). Here, the parameters $\gamma,\gamma_{e}$ define the strength of the gain modulation, as explained in the Methods section. These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$. We refer to Supporting Information for further details on the training procedure.
In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN. This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 2B).
@@
In Fig.2B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line.
The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.2C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.2C).
@@
Figure 3: Dynamical reinforcement learning: Multi Armed Bandits. A. Schematic of network architecture and task. Virtual weighs parameterize the agent’s policy (orange). At each round, an action (red) is sampled from the policy and played. A reward is then sampled from the current environment. A GM-network (green) then predicts the virtual parameter update. B. Comparing in distribution (ID) and out-of-distribution (OOD) expected reward at the 1000-th round for the policy gradient (PG) training source, reward gain-modulated network ($\gamma>0$), and a network without gain modulation $\gamma=0$. C. ID and OOD model performance varying gain modulation strength ($\gamma$). Solid lines indicate the median score (expected reward), computed over 20 independently trained models. The filled area indicates 25-75% confidence interval. D. Same as B but comparing OOD performance for different network sizes (number of hidden units). E. Robustness to hyperparameter variations in architectures with $(\gamma=0.9)$ and without ($\gamma=0.0$) gain modulation. We report the percentage of models that fail to generalize to OOD environments, varying the standard deviation of apical ($\sigma_{a}$) and basal ($\sigma_{b}$) random input weights. F. Regret comparison between the distilled and the original policy gradient algorithm. We report average regret per round in ID and OOD setting, compared with the one achieved by the PG training source.
Multi Armed Bandits
@@
Dynamical adaptation is now investigated within the framework of reinforcement learning. We provide robust evidence supporting the hypothesis that gain modulation represents a fundamental component in implementing in-context learning in biological agents. We first explore stateless environments $\mathcal{A}^{bandits}$, which are represented by Bernoulli K-armed bandits [32]. Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability. During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
@@
In this simplified bandit scenario, where state information is absent, the virtual weights $\bm{w}$ directly parameterize the policy probabilities: $\bm{\pi}=\mathrm{softmax}{(\bm{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 3A, green). This serves as an ideal test bed for evaluating the network ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action, as described in the Methods section. We systematically investigate the impact of the gain modulation on out-of-distribution performance, varying the parameter $\gamma$. Notably, when $\gamma=0$ the reward does not modulate the network gain.
We present our experimental findings in Fig. 3. Initially, we observe that integrating reward gain modulation into the network enables the agent to consistently converge to the optimal policy, even in out-of-distribution environments (see Fig.3B). To further explore this phenomenon, we trained multiple models while varying the strength of the modulation parameter ($\gamma$). We observe that performance improves with higher values of $\gamma$, allowing the generalization of learned rules out-of-distribution (Fig. 3C) and showcasing increased robustness to variations in hyperparameters (Fig. 3E). Furthermore, we observe that in the absence of gain modulation, the network is unable to generalize, merely memorizing the rules from the training set. This is evidenced by its persistence in selecting even-numbered arms even when tested in out-of-distribution environments where high rewards are allocated to odd-numbered arms (see Fig. 3E). Comparing the performances of models of different sizes (number $N$ of hidden units), we see that models with more reward gain modulation need significantly fewer neurons to achieve maximum scores in OOD environments (see Fig.3D).
Analysing the regret curves displayed in Fig. 3F we observe that a model with gain modulation learns a more data-efficient algorithm than its source even in OOD environments. On the contrary, a model without gain modulation fails to generalize to OOD environments, not converging to zero regret (see Fig. 3F, right panel).
@@
Figure 4: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach unobservable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). E. Reward plotted against the number of trials, for the policy gradient (PG) algorithm (displayed in blue lines) applied in the dynamical PG scenario on both training positions (depicted in the left panel) and new positions (shown in the right panel). Thin line: single position, thick line: average across different positions. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
Reinforcement learning in a reaching task.
@@
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [33]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility.
The horizontal and vertical positions of the agent are encoded separately using five input units $\bm{x}$ each, employing Gaussian activation functions with equally spaced centers ($0.4$ apart) and a width of $0.4$.
@@
To achieve this task we consider a network architecture composed of two networks. One GM-network, , $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 4A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig. 4A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equations:
$\begin{cases}\bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq\bm{x}\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\end{cases}$ (11)
Indeed, we compare the case with (Fig. 4A) and without (Fig. 4B) gain modulation. We refer to Supporting Information for further details on the training procedure.
@@
To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig. 4C, left panel) and new test positions (OOD, Fig. 4C, right panel). For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training. The agent’s precision closely matches that of the plastic policy gradient learning rule. We present the reward plotted against the number of trials for both training (Fig. 4E, right panel, pink lines) and test (Fig. 4C, right panel, pink lines) food locations.
@@
We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig. 4B, D, F). Notably, while performances for ID food locations are acceptable (Fig. 4E, left panel, pink lines), those for OOD cases are extremely low (Fig. 4E, right panel, pink lines). This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
Discussion
@@
It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 34], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8]. Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers. One intriguing explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11]. However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals. In this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
@@
When considering a stationary dataset our model dynamically adapt to the context provided by the data, allowing predictions on new data, without synaptic plasticity. Remarkably, when the learning rule shifts abruptly, the network swiftly adapts without retraining, mimicking human cognitive flexibility observed in tasks like the Wisconsin Card Sorting Test. Conversely, when using gradient-based plasticity, the learning speed resembles what observed in certain pathological conditions [31]. When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, that are updated in function of the reward. We stress that virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other RNNs, that were pre-trained to perform gradient updates. Another pre-trained network receives those weights as an input, along with the current environment features, and computes an adapted in-context response. We find that networks with gain modulation exhibit improved performance and robustness to variations compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
@@
In conclusion, our approach provide an explicit framework to induce ICL in biologically plausible network, possibly opening the route to a formal understanding of ICL in biological agents.
Acknowledgments

### v2

@@ 3 Results
Figure 1: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of $N=20$ units (described by the vector $\bm{z}(t)$, depicted in pink, which follows Eq. (6)) is utilized for learning a periodic trajectory, following the prescription of reservoir computing. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights $\delta w$, a $N$-dimensional vector (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (solid: median, dashed: 20-th/80-th percentile (statistics evaluated over 10 realizations).
@@
We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm, we employ an RNN (Fig. 1A, pink), whose dynamics follows Equation (6), to extract temporal features. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x=y^{targ}(t)$ in Eq. (6), while $R$ is Gaussian matrix with zero mean and variance, $\sigma_{R}^{2}$) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 1B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN. The readout virtual weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
@@
Here, the network $W_{\Theta^{w}}$$({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)=\textrm{GMNN}_{\Theta^{w}}($$\bm{z}$$,e|\gamma_{e}e)$ is dedicated to estimating the gradient of virtual weights, while $Y_{\Theta^{y}}$(((($\bm{z}$$,w)=\textrm{GMNN}_{\Theta^{y}}($$\bm{z}$$,w|\gamma w)$ is tasked with estimating the predicted $y(t)$ as a function of the new virtual weights (Fig.1A green and orange respectively). Here, the parameters $\gamma,\gamma_{e}$ define the strength of the gain modulation, as explained in the Methods section.
These gain-modulated architectures are pre-trained to replicate, respectively, gradient descent updates and scalar products obtained on a set of $N_{train}$ training sequences $\{\{y_{\alpha}^{targ}(t)\}_{t}:\ \alpha\in[N_{train}]\}$. More specifically, the target sequences are $5$ sinusoidal functions with different frequencies (see vertical red lines in Fig. 1C.)
@@
In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq. (6)). This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 1B). In Fig.1B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line. The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.1C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.1C, solid and dashed lines represent respectively, median and 20th/80th percentile range. Statistics is evaluated over 10 realizations of the experiment).
@@
Figure 2: Dynamical reinforcement learning: Multi Armed Bandits. A. Schematic of network architecture and task. Virtual weighs parameterize the agent’s policy (orange). At each round, an action (red) is sampled from the policy and played. A reward is then sampled from the current environment. A GM-network (green) then predicts the virtual parameter update. B. Regret comparison between the distilled and the original policy gradient algorithm. We report log regret per round distribution achieved at the 100-th round for 100 independently trained models in ID and OOD settings. We compare a gain-modulated network ($\gamma=1$), and a network without gain modulation ($\gamma=0$) with the policy gradient training source (PG). C. OOD model performance varying number of hidden dimensions. Solid lines indicate the median score (expected reward), computed over 100 independently trained models. The filled area indicates 20-80% confidence interval. D. We report average regret per round curves in OOD setting. We compare a gain-modulated network ($\gamma=1$, teal), and a network without gain modulation ($\gamma=0$, orchid) with the policy gradient training source (in black). Each curve represents the average regret per round for one fixed trained model, computed by averaging over 100 independent simulations.
3.1 Multi Armed Bandits
@@
Dynamical adaptation is now investigated within the framework of reinforcement learning. We provide robust evidence supporting the hypothesis that gain modulation represents a fundamental component in implementing in-context learning in biological agents. We first explore stateless environments $\mathcal{A}^{bandits}$, which are represented by Bernoulli K-armed bandits [30]. Within each environment $\alpha\in\mathcal{A}^{bandits}$, there exists a subset $P_{\alpha}\subset[K]$ of arms that yield a reward with high probability ($p=0.95$). During the training phase, the in-distribution environments give a high reward probability to even-numbered arms, whereas during testing, the out-of-distribution environments assign a high reward probability to odd-numbered arms.
@@
In this simplified bandit scenario, where state information is absent, the virtual weights $\bm{w}$ directly parameterize the policy probabilities: $\bm{\pi}=\mathrm{softmax}{(\bm{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 2A, green). This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action. We train this network to approximate the policy gradient update rule. As training data, we use policy gradient estimates over one single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment. We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) models in out-of-distribution (OOD) environments. We train the network to approximate the policy gradient update rule using policy gradient estimates from a single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment. We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) in out-of-distribution (OOD) environments.
To systematically investigate the impact of gain modulation ($\gamma=1$) on out-of-distribution performance, we compare it with networks where the reward does not modulate the network gain ($\gamma=0$). For each case, we select the optimal hyperparameters through a grid-based search (additional details in the Appendix).
Our experimental findings are presented in Fig.2. First, comparing models of different sizes (number $N$ of hidden units), we find that models with gain modulation require significantly fewer neurons to achieve maximum scores in OOD environments (Fig.2C).
In Fig.2B, we report the regret per round distribution at the 100th round and compare it with the distribution obtained by policy gradient with the same learning rate and iterations. A gain-modulated network achieves a regret distribution comparable to the policy gradient in both ID and OOD environments, with a lower median. In contrast, networks without gain modulation show significantly higher regret. Analyzing the regret curves in Fig.2D, we observe that a model with gain modulation often learns a more data-efficient algorithm than its source, even in OOD environments. Conversely, a model without gain modulation fails to generalize to OOD environments and does not converge to zero regret.
In summary, gain modulation enables the network to consistently and efficiently distill the correct gradient update rule and generalize it to unseen environments, predicting the correct virtual weight update in regions of the input space far from the training data.
@@
Figure 3: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). C. Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD). Line: median, shading: 20-th/80-th percentile range. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
3.2 Reinforcement learning in a reaching task.
@@
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [31]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility. The position of the agent is encoded separately using 25 input units $\bm{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
@@
To accomplish this task we consider a network architecture composed of two networks. One GM-network, $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 3A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig. 3A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equation Eq.(5). Indeed, we compare the case with (Fig. 3A) and without (Fig. 3B) gain modulation. We refer to Supporting Information for further details on the training procedure.
@@
To validate that our trained model is capable of dynamically implementing policy gradient itself, we tested it on both the training set locations (ID, Fig. 3C, left panel) and new test positions (OOD, Fig. 3C, right panel). For each food position (coded with different colors), we illustrate a sample trajectory executed by the agent (in corresponding colors) to reach the target at the end of the training. The agent’s precision closely matches that of the plastic policy gradient learning rule. We present the reward plotted against the number of trials for both training (Fig. 3E, right panel, pink line) and test (Fig. 3C, right panel, pink line) food locations.
@@
We compared these performances against an architecture lacking gain modulation ($\gamma=0$), observing worst performances (see Fig. 3B, D, F). Notably, while performances for ID food locations are acceptable (Fig. 3E, left panel, pink lines), those for OOD cases are extremely poor (Fig. 3E, right panel, pink lines). This observation, coupled with the results from the preceding section, suggests that gain modulation is a crucial component in facilitating the generalization of adaptive behavioral capabilities in recurrent networks.
We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section C in Appendix), requiring evaluating temporal credit assignment.
4 Discussion
@@
It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 32], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8]. Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers. This explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11]. However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals. In this this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
@@
When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, that are updated in function of the reward. We stress that virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other RNNs, that were pre-trained to perform gradient updates. Another pre-trained network receives those weights as an input, along with the current environment features, and computes an adapted in-context response. We find that networks with gain modulation exhibit improved performance and robustness to variations compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
@@
In conclusion, our approach provides an explicit framework to induce ICL in biologically plausible networks, possibly opening the route to a formal understanding of ICL in biological agents.
Limitations of this study
In this study, we focused on dynamical adaptation of virtual readout weights, in other words, only the last layer is considered. We are neglecting the dynamics on recurrent weights and weights of the hidden units, which, in general, strongly affect performance and stability. Generalizability to untested real-world conditions is uncertain, and we do not investigate the scalability to larger architectures. We discuss that the mechanism we discuss could be present in biological agents, allowing in-context learning. However, the theoretical basis for attention mechanisms executing gradient descent updates lacks comprehensive empirical validation in biological experiments.
Additionally, the study does not fully address how dynamic gradient structures could be learned or developed in biological networks.
5 Acknowledgments